Find the Polynoial of lowest degree that has the zeros, 4,-4, 1, -1.

Answer by rapaljer(4671)

P(X)= (x-4)(x+4)(x-1)(x+1)
P(x) = (x^2-16)(x^-1)
P(x)=x^4 -17x^2+16

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus